# Generalized and Unified Equivalences between Hardness and Pseudoentropy

### v1

@@ 4 Generalized and Unified Pseudoentropy Characterizations
In this section, we complete the proofs of our main results Theorems 1.10 and 1.11 using the enhanced regularity lemmas Theorems 3.1 and 3.10 from the previous section.
@@ 4.1 Non-uniform Setting
Now the first half of Theorem 1.10 (up to Equation (4)) follows directly from Theorem 3.1, where we set :=$\mathcal{R}:=\{r_{\varphi}\}_{\varphi\in\Phi},T_{\mathcal{R}}:=T_{\Phi}$, set $\mathcal{F}$ to be the family of functions $f:\{0,1\}^{n}\to[-1,1]^{L}$ with circuit complexity at most $T$, set $T_{\mathcal{F}}:=T$, and replace $\varepsilon$ with $\varepsilon/2$.
To prove the second half of Theorem 1.10 (i.e. the reverse direction), we note the following identity for every $g,g^{*},s:\{0,1\}^{n}\to\Delta_{L}$:
@@
This identity follows immediately from the definitions of $\mathsf{H}_{\varphi}$ and $\mathsf{D}_{\varphi}$ in Definitions 1.7 and 1.8. By (37) and the fact that $\mathsf{D}_{\varphi}(s\|g)\geq 0$, we have
@@ 4.2 Uniform Setting
We need the following definition to give a formal statement of Theorem 1.11.
@@
The following claim follows immediately from Definitions 3.9 and 4.1:
@@
Theorem 4.3 (Formal statement of Theorem 1.11).
@@
    If in addition $\mathcal{B}$ is an ($(\{\nabla\varphi\}_{\varphi\in\Phi},\varepsilon_{1},\varepsilon,\delta^{\prime})$ weak agnostic calibration oracle for some $\varepsilon_{1}>0$, then with probability at least $1-\delta$ over the random draw of the input data points, the output function $s$ is $(\mathcal{A},\varepsilon,\delta^{\prime})$-indistinguishable from $g^{*}$ and satisfies
@@
The theorem now follows directly from Theorem 3.10 and 4.2, where we choose =$\mathcal{R}=\{\nabla\varphi\}_{\varphi\in\Phi}$. ∎
@@ 5 Handling Unbounded Subgradients
Our Theorem 1.10 requires the assumption that the subgradient $\nabla\varphi$ is bounded. This assumption is not satisfied by the Shannon entropy, where $\varphi(v)=\sum v_{i}\ln v_{i}$. In this case, $\nabla\varphi(v)=(\ln(v_{1}),\ldots,\ln(v_{L}))$ becomes unbounded when some $v_{i}$ approaches zero. We show that this can be easily fixed using the following more general theorem than Theorem 1.10. Here, we consider an arbitrary transformation $\sigma_{\varphi}:\Delta_{L}\to\Delta_{L}$ associated with each $\varphi\in\Phi$, and Theorem 1.10 is the special case when these transformations $\sigma_{\varphi}$ are the identity function. We will later choose these transformations so that the subgradient $\nabla\varphi(\sigma_{\varphi}(v))$ at the transformed point $\sigma_{\varphi}(v)$ is bounded, even when $\nabla\varphi(v)$ may not be bounded.
@@
Let $T>0$ be a time bound and let $\varepsilon\in(0,1)$ be an error parameter. Let $\Phi$ be a family of convex functions $\varphi:\Delta_{L}\to\mathbb{R}$. For every $\varphi\in\Phi$, let $\sigma_{\varphi}:\Delta_{L}\to\Delta_{L}$ be an arbitrary transformation. Assume that given $v\in\Delta_{L}$, the transformed subgradient $\nabla\varphi(\sigma_{\varphi}(v))$ is bounded in $[-1,1]^{L}$ and can be computed to $\ell_{\infty}$ accuracy $\varepsilon/4$ in nonuniform time $T_{\Phi}$, for every $\varphi\in\Phi$. Then there exists
@@
Theorem 5.1 can be proved in the exact same way as Theorem 1.10 using the enhanced regularity lemma Theorem 3.1, and similarly, we can prove a generalized version of Theorem 4.3 using Theorem 3.10 to handle unbounded subgradients in the uniform setting.
The advantage of using Theorem 5.1 is that it only requires $\nabla\varphi(\sigma_{\varphi}(v))$ to be bounded after the transformation $v\mapsto\sigma_{\varphi}(v)$. The challenge, however, is that while (41) is stated for the transformed $\sigma_{\varphi}\circ s$, Theorem 5.1 only ensures that the un-transformed $s$ belongs to $\mathsf{Ind}(g^{*};T,\varepsilon)\cap\mathsf{Time}(T^{\prime})$. This mis-alignment between $s$ and $\sigma_{\varphi}\circ s$ also appears in the reverse direction (42) where we minimize the divergence of the transformed $\sigma_{\varphi}\circ g$ subject to the un-transformed $g$ being in $\mathsf{Time}(T)$. Below we show how to address this mis-alignment challenge for the Shannon entropy by choosing the transformation $\sigma_{\varphi}$ appropriately.
For the (negative) Shannon entropy $\varphi$ in (25), define $\sigma_{\varphi}(v)=(1-\varepsilon)v+\varepsilon u$, where $u=(1/L,\ldots,1/L)$ is the uniform distribution and $\varepsilon\in(0,1/2)$. Intuitively, this transformation perturbs $v$ away from the boundary of $\Delta_{L}$. We now have $\|\nabla\varphi(\sigma_{\varphi}(v))\|_{\infty}\leq\ln(L/\varepsilon)$, so the boundedness assumption of Theorem 5.1 is satisfied after a multiplicative scaling of $1/\ln(L/\varepsilon)$. Moreover, we have the following implication connecting $\sigma_{\varphi}\circ s$ with $s$ that addresses the mis-alignment challenge:
@@ 6 Exponential Lower Bound for a Particular Entropy Notion
We prove Theorem 1.12 and Corollary 1.13 in this section. We need the following helper lemma:
@@
We define $\alpha:=1/16$. By Lemma 6.1, there exists a subset $\mathcal{X}^{\prime}\subseteq\{0,1\}^{n}$ such that
@@
$\nabla\varphi(v)$ $:=x_{v}^{(1-\eta(x_{v}))}\in\{0,1\}^{L},\quad\text{where $x_{v}=\operatorname*{arg\,max}_{x\in\mathcal{X}^{\prime}}\langle v,x^{(1-\eta(x))}\rangle$}.$
@@
$-\mathsf{H}_{\varphi}(g^{*})\leq 4\alpha.$ (43)
@@
$-\mathsf{H}_{\varphi}(g^{*})=\mathbb{E}_{x^{\prime}\sim\mu}\left[\max_{x\in\mathcal{X}^{\prime}}\langle g^{*}(x^{\prime}),x^{(1-\eta(x))}\rangle\right].$ (44)
@@
Now let $s:\{0,1\}^{n}\to\Delta_{L}$ be an arbitrary function that is ($(\{f\},0.05)$ indistinguishable from $g^{*}$ and satisfies (7). By the indistinguishability, we have
@@
$\mathbb{E}_{x\sim\mu}\langle s(x),x^{(1-\eta(x))}\rangle=-\mathsf{H}_{\varphi}(s)\leq-\mathsf{H}_{\varphi}(g^{*})+0.05\leq 4\alpha+0.05=1/4+0.05.$
@@
Proof of Corollary 1.13.
It suffices to show that the two conditions (multiaccuracy and weight-restricted calibration) satisfied by $s$ imply that $s$ belongs to $\mathsf{Ind}(g^{*};\{f\},0.05)$ and satisfies (7) as required by Theorem 1.12.
@@ Appendix A Deferred Proofs
[…]
@@
[…]
@@
$\Gamma(g,h_{0})$ $\geq\frac{1}{2\lambda}\|g-g_{0}\|^{2}$ for every $\text{for every $g\in\mathcal{G}$},$
[…]
@@
$\Gamma(g_{0},h_{0}+h)$ $\leq\frac{\lambda}{2}\|h\|_{*}^{2}$ for every $\text{for every $h\in\mathbb{R}^{L}$},$
[…]
@@
This shows that $\varphi$ is $1$-strongly convex w.r.t. $\|\cdot\|_{1}$. By A.3, we know that $\psi$ is $1$-smooth w.r.t. $\|\cdot\|_{\infty}$. Let $(g,h),(\hat{g},\hat{h})\in\mathcal{G}\times\mathbb{R}^{L}$ be two arbitrary conjugate pairs. By (48) and the $1$-strong convexity of $\varphi$ w.r.t. $\|\cdot\|_{1}$, we have
@@
Since this holds for arbitrary $h,\hat{h}\in\mathbb{R}^{L}$ and the gradients $g:=\nabla\psi(h)=\mathsf{softmax}(h),\hat{g}:=\nabla\psi(\hat{h})=\mathsf{softmax}(\hat{h})$, we have proved the $1$-Lipschitzness of softmax. ∎

### v2

@@ 4 Generalized and Unified Pseudoentropy Characterizations
In this section, we complete the proofs of our main results Theorems˜1.10 and 1.11 using the enhanced regularity lemmas Theorems˜3.1 and 3.10 from the previous section.
@@ 4.1 Non-uniform Setting
Now the first half of Theorem˜1.10 (up to Equation (4)) follows directly from Theorem˜3.1, where we set $\mathcal{R}:=\{r_{\varphi}\}_{\varphi\in\Phi},T_{\mathcal{R}}:=T_{\Phi}$, set $\mathcal{F}$ to be the family of functions $f:\{0,1\}^{n}\to[-1,1]^{L}$ with circuit complexity at most $T$, set $T_{\mathcal{F}}:=T$, and replace $\varepsilon$ with $\varepsilon/2$.
To prove the second half of Theorem˜1.10 (i.e. the reverse direction), we note the following identity for every $g,g^{*},s:\{0,1\}^{n}\to\Delta_{L}$:
@@
This identity follows immediately from the definitions of $\mathsf{H}_{\varphi}$ and $\mathsf{D}_{\varphi}$ in Definitions˜1.7 and 1.8. By (37) and the fact that $\mathsf{D}_{\varphi}(s\|g)\geq 0$, we have
@@ 4.2 Uniform Setting
We need the following definition to give a formal statement of Theorem˜1.11.
@@
The following claim follows immediately from Definitions˜3.9 and 4.1:
@@
Theorem 4.3 (Formal statement of Theorem˜1.11).
@@
    If in addition $\mathcal{B}$ is an $(\{\nabla\varphi\}_{\varphi\in\Phi},\varepsilon_{1},\varepsilon,\delta^{\prime})$ weak agnostic calibration oracle for some $\varepsilon_{1}>0$, then with probability at least $1-\delta$ over the random draw of the input data points, the output function $s$ is $(\mathcal{A},\varepsilon,\delta^{\prime})$-indistinguishable from $g^{*}$ and satisfies
@@
The theorem now follows directly from Theorem˜3.10 and ˜4.2, where we choose $\mathcal{R}=\{\nabla\varphi\}_{\varphi\in\Phi}$. ∎
@@ 5 Handling Unbounded Subgradients
Our Theorem˜1.10 requires the assumption that the subgradient $\nabla\varphi$ is bounded. This assumption is not satisfied by the Shannon entropy, where $\varphi(v)=\sum v_{i}\ln v_{i}$. In this case, $\nabla\varphi(v)=(\ln(v_{1}),\ldots,\ln(v_{L}))$ becomes unbounded when some $v_{i}$ approaches zero. We show that this can be easily fixed using the following more general theorem than Theorem˜1.10. Here, we consider an arbitrary transformation $\sigma_{\varphi}:\Delta_{L}\to\Delta_{L}$ associated with each $\varphi\in\Phi$, and Theorem˜1.10 is the special case when these transformations $\sigma_{\varphi}$ are the identity function. We will later choose these transformations so that the subgradient $\nabla\varphi(\sigma_{\varphi}(v))$ at the transformed point $\sigma_{\varphi}(v)$ is bounded, even when $\nabla\varphi(v)$ may not be bounded.
@@
Let $T>0$ be a time bound and let $\varepsilon\in(0,1)$ be an error parameter. Let $\Phi$ be a family of convex functions $\varphi:\Delta_{L}\to\mathbb{R}$. For every $\varphi\in\Phi$, let $\sigma_{\varphi}:\Delta_{L}\to\Delta_{L}$ be an arbitrary transformation. Assume that given $v\in\Delta_{L}$, the transformed subgradient $\nabla\varphi(\sigma_{\varphi}(v))$ is bounded in $1,1]^{L}$ and can be computed to $\ell_{\infty}$ accuracy $\varepsilon/4$ in nonuniform time $T_{\Phi}$, for every $\varphi\in\Phi$. Then there exists
@@
Theorem˜5.1 can be proved in the exact same way as Theorem˜1.10 using the enhanced regularity lemma Theorem˜3.1, and similarly, we can prove a generalized version of Theorem˜4.3 using Theorem˜3.10 to handle unbounded subgradients in the uniform setting.
The advantage of using Theorem˜5.1 is that it only requires $\nabla\varphi(\sigma_{\varphi}(v))$ to be bounded after the transformation $v\mapsto\sigma_{\varphi}(v)$. The challenge, however, is that while (41) is stated for the transformed $\sigma_{\varphi}\circ s$, Theorem˜5.1 only ensures that the un-transformed $s$ belongs to $\mathsf{Ind}(g^{*};T,\varepsilon)\cap\mathsf{Time}(T^{\prime})$. This mis-alignment between $s$ and $\sigma_{\varphi}\circ s$ also appears in the reverse direction (42) where we minimize the divergence of the transformed $\sigma_{\varphi}\circ g$ subject to the un-transformed $g$ being in $\mathsf{Time}(T)$. Below we show how to address this mis-alignment challenge for the Shannon entropy by choosing the transformation $\sigma_{\varphi}$ appropriately.
For the (negative) Shannon entropy $\varphi$ in (25), define $\sigma_{\varphi}(v)=(1-\varepsilon)v+\varepsilon u$, where $u=(1/L,\ldots,1/L)$ is the uniform distribution and $\varepsilon\in(0,1/2)$. Intuitively, this transformation perturbs $v$ away from the boundary of $\Delta_{L}$. We now have $\|\nabla\varphi(\sigma_{\varphi}(v))\|_{\infty}\leq\ln(L/\varepsilon)$, so the boundedness assumption of Theorem˜5.1 is satisfied after a multiplicative scaling of $1/\ln(L/\varepsilon)$. Moreover, we have the following implication connecting $\sigma_{\varphi}\circ s$ with $s$ that addresses the mis-alignment challenge:
@@ 6 Exponential Lower Bound for a Particular Entropy Notion
We prove Theorem˜1.12 and Corollary˜1.13 in this section. We need the following helper lemma:
@@
We define $\alpha:=1/16$. By Lemma˜6.1, there exists a subset $\mathcal{X}^{\prime}\subseteq\{0,1\}^{n}$ such that
@@
[…]
@@
$\mathsf{H}_{\varphi}(g^{*})\leq 4\alpha.$ (43)
@@
$\mathsf{H}_{\varphi}(g^{*})=\mathbb{E}_{x^{\prime}\sim\mu}\left[\max_{x\in\mathcal{X}^{\prime}}\langle g^{*}(x^{\prime}),x^{(1-\eta(x))}\rangle\right].$ (44)
@@
Now let $s:\{0,1\}^{n}\to\Delta_{L}$ be an arbitrary function that is $(\{f\},0.05)$ indistinguishable from $g^{*}$ and satisfies (7). By the indistinguishability, we have
@@
$\mathbb{E}_{x\sim\mu}\langle s(x),x^{(1-\eta(x))}\rangle\leq-\mathsf{H}_{\varphi}(s)\leq-\mathsf{H}_{\varphi}(g^{*})+0.05\leq 4\alpha+0.05=1/4+0.05.$
@@
Proof of Corollary˜1.13.
It suffices to show that the two conditions (multiaccuracy and weight-restricted calibration) satisfied by $s$ imply that $s$ belongs to $\mathsf{Ind}(g^{*};\{f\},0.05)$ and satisfies (7) as required by Theorem˜1.12.
@@ Appendix A Deferred Proofs
Let $\mathcal{G}\subseteq\mathbb{R}^{L}$ be a bounded convex set. For $\lambda\geq 0$ and a norm $\|\cdot\|$ on $\mathbb{R}^{L}$, we say a convex function $\varphi:\mathcal{G}\to\mathbb{R}$ is $\lambda$-strongly convex w.r.t. $\|\cdot\|$ if for every $g_{0},g\in\mathcal{G}$, letting $h_{0}\in\mathbb{R}^{L}$ be an arbitrary subgradient of $\varphi$ at $g_{0}$, we have
@@
Let $\mathcal{G}\subseteq\mathbb{R}^{L}$ be a bounded convex set. For $\lambda>0$, if $\varphi:\mathcal{G}\to\mathbb{R}$ is a closed $\frac{1}{\lambda}$-strongly convex function w.r.t a norm $\|\cdot\|$ on $\mathbb{R}^{L}$, then its convex conjugate $\psi$ is $\lambda$-smooth w.r.t. the dual norm $\|\cdot\|_{*}$.
@@
$\Gamma(g,h_{0})$ $\geq\frac{1}{2\lambda}\|g-g_{0}\|^{2}$ $\text{for every $g\in\mathcal{G}$},$
$\text{or equivalently,}\quad\Gamma(g+g_{0},h_{0})$ $\geq\frac{1}{2\lambda}\|g\|^{2}$ $\text{for every $g$ satisfying $g+g_{0}\in\mathcal{G}$}.$
@@
$\Gamma(g_{0},h_{0}+h)$ $\leq\frac{\lambda}{2}\|h\|_{*}^{2}$ $\text{for every $h\in\mathbb{R}^{L}$},$
$\hskip 50.00008pt\text{or equivalently,}\quad\Gamma(g_{0},h)$ $\leq\frac{\lambda}{2}\|h-h_{0}\|_{*}^{2}$ $\text{for every $h\in\mathbb{R}^{L}$}.\hskip 50.00008pt$
@@
This shows that $\varphi$ is $1$-strongly convex w.r.t. $\|\cdot\|_{1}$. By ˜A.3, we know that $\psi$ is $1$-smooth w.r.t. $\|\cdot\|_{\infty}$. Let $(g,h),(\hat{g},\hat{h})\in\mathcal{G}\times\mathbb{R}^{L}$ be two arbitrary conjugate pairs. By (48) and the $1$-strong convexity of $\varphi$ w.r.t. $\|\cdot\|_{1}$, we have
@@
[…]